OpenStudy (anonymous):

in a recent year, the total sales at Restaurant A were $25,000 more than at Restaurant B. Total sales at the 2 stores were $600,000,000. What were the sales for the year at each store?

OpenStudy (anonymous):

A=25000 + B A+B=600000000 REPLACE A IN THE SECOND EQUATION 25000 +B +B =600000000 2B =599975000 B=299987500 A=25000+B A= 300012500